OpenStudy (anonymous):

An architect is considering bidding for the design of a new shopping mall. The cost of drawing plans and submitting a model is $10,000. The probability of being awarded the bid is 0.11, and anticipated profits are $100,000, resulting in a possible gain of this amount minus the $10,000 cost for plans and a model. What is the expected value in this situation?

OpenStudy (anonymous):

if x=10,000 the probability 1-0.11=0.89 x=100,000-10,000=90,000 , prob=0.11 expected value E=90,000*0.11 -10,000*0.89 =1000 dollar or try E=90,000*.89 - 10,000*0.11=79,000 dollar i think its the lower one 1000 dollar